roxy’s ice cream shop sells ice cream sandwiches for $2 and root beer floats for $3. on a hot day the shop sold 8 more sandwiches than root beet floats and made $156 how many root beer floats did they sell?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52788)

.

Let x be the # of root beer floats.


Then the # of sandwiches is (x+8).


Next, you write the total money equation


    2*(x+8) + 3x = 156   dollars.


From the equation


    2x + 16 + 3x = 156

    5x          = 156 - 16 =  140

     x                     = 140/5 = 28  root beer floats sold.

Solved.




Answer by greenestamps(13200)



The "extra" 8 sandwiches cost 8($2) = $16.

The remaining items are equal numbers of sandwiches and floats, for a total of $156-$16 = $140.

The cost of one sandwich and one float is $2+$3 = $5.

The number of sandwiches and floats that must be sold to make $140 is $140/$5 = 28.

So the number of root beer floats sold was 28. (And the number of sandwiches sold was 28+8 = 36; but the problem didn't ask for that....).

Now compare this informal solution to the formal algebraic solution shown by the other tutor. You will see that the calculations are virtually the same.

That shows you that algebra is not "magic" -- it is just a formalization of what makes logical sense.

So don't be afraid of the algebra; learn how to use it. When the problems get more complicated, an informal solution might be difficult or impossible, and you will need the algebra.